OpenStudy (firejay5):

Show work!!!!!!!!! 9. A museum curator moves artifacts into place on various different display surfaces. If the curator moves a 145kg aluminum sculpture across a horizontal steel platform with a force of 668N, what is the coefficient of kinetic friction?

OpenStudy (firejay5):

My Answer: Fn = m*g Fn = (145) (9.8) Fn = 1,421N Uk = Fk/Fn ; Uk = 668N/1,421N ; Uk = 0.47

OpenStudy (firejay5):

@midhun.madhu1987

OpenStudy (firejay5):

Is my answer correct

OpenStudy (midhun.madhu1987):

yes... i think so

OpenStudy (firejay5):

I have two more questions for help real quick

OpenStudy (midhun.madhu1987):

I am not good at Physics...i'll try.. tag me in the question